# Learning Grouped Lattice Vector Quantizers for Low-Bit LLM Compression

### v1

@@ 4 Experiments
In this section, we evaluate the performance of our proposed grouped Lattice vector quantization (GLVQ) framework on the Llama 1 and 2 models touvron2023llama , which range from 7 billion to 70 billion parameters. To ensure a fair comparison, we follow the experimental setup used in QuIP#. Specifically, we evaluate perplexity on the Wikitext-2 wikitext2 and C4 c4 datasets, utilizing context lengths of 2048 for Llama 1 and 4096 for Llama 2 models. For zero-shot tasks, we use the LM Eval framework to measure accuracy on tasks such as ARC, PIQA, and the Winograd Schema Challenge (Wino). We report both perplexity and zero-shot accuracy for all methods to offer a comprehensive evaluation. We implement our method using PyTorch pytorch and CUDA cuda , with all experiments conducted on NVIDIA A100 GPUs. For timing experiments, we use an NVIDIA RTX 4090 GPU. We adopt 4M tokens from the RedPajama 1T dataset together2023redpajama as the calibration sequences in our experiments. We implement two variants of our model with lattice dimensions $d=8$ and $d=32$, referred to as GLVQ-8D and GLVQ-32D, respectively. For baselines, we compare our method against several state-of-the-art PTQ approaches, including OmniQuant shao2023omniquant , AWQ lin2024awq QuIP# tseng2024quip , AQLM egiazarian2024extreme and QTIP tseng2024qtip .
@@ 4.1 Perplexity Results on Llama Models
Our evaluation focuses on perplexity over the Wikitext-2 wikitext2 and C4 c4 datasets, using a context length of 2048 for Llama 1 and 4096 for Llama 2. Results are reported in Table 1. Across all settings, GLVQ consistently achieves lower perplexity than existing methods, particularly in low-bit regimes. For example, under the 2-bit quantization regime on Llama 2-70B, GLVQ-32D achieves a perplexity of 3.36 on Wikitext-2, significantly lower than QTIP (3.78) and QuIP# (3.91), highlighting that the performance gap becomes most pronounced under extreme compression levels. We also observe consistent improvements across model scales and datasets. On Llama 1-13B, GLVQ-32D improves upon QuIP# by 0.41 perplexity points (5.38 vs. 5.79) at 2-bit on Wikitext-2, and 0.53 points (6.95 vs. 7.48) on C4. These results validate the robustness of our method across both small and large Llama variants. Notably, the performance gain is more pronounced as the bit-width decreases, highlighting the advantage of GLVQ’s group-specific lattice codebooks and companding functions. The combination of adaptive vector quantization and nonlinear transformation helps mitigate the quantization error that typically worsens at extreme compression levels. Furthermore, GLVQ-32D consistently outperforms GLVQ-8D, demonstrating that larger lattice dimensions improve quantization fidelity by offering a more expressive codebook structure. This supports our design choice of learning group-specific lattices with adjustable resolution.

### v2

@@ 4 Experiments
In this section, we evaluate the performance of our proposed grouped Lattice vector quantization (GLVQ) framework on the Llama 1 and 2 models Touvron et al. (2023), which range from 7 billion to 70 billion parameters. To ensure a fair comparison, we follow the experimental setup used in QuIP#. Specifically, we evaluate perplexity on the Wikitext-2 Merity et al. (2016) and C4 Raffel et al. (2020) datasets, utilizing context lengths of 2048 for Llama 1 and 4096 for Llama 2 models. For zero-shot tasks, we use the LM Eval framework to measure accuracy on tasks such as ARC, PIQA, and the Winograd Schema Challenge (Wino). We report both perplexity and zero-shot accuracy for all methods to offer a comprehensive evaluation. We implement our method using PyTorch Paszke et al. (2017) and CUDA NVIDIA (2020), with all experiments conducted on NVIDIA A100 GPUs. For timing experiments, we use an NVIDIA RTX 4090 GPU. We adopt 4M tokens from the RedPajama 1T dataset Weber et al. (2024) as the calibration sequences in our experiments. We implement two variants of our model with lattice dimensions $d=8$ and $d=32$, referred to as GLVQ-8D and GLVQ-32D, respectively. For baselines, we compare our method against several state-of-the-art PTQ approaches, including OmniQuant Shao et al. (2023), AWQ Lin et al. (2024) QuIP# Tseng et al. (2024a), AQLM Egiazarian et al. (2024) and QTIP Tseng et al. (2024b).
@@ 4.1 Perplexity Results on Llama Models
Our evaluation focuses on perplexity over the Wikitext-2 Merity et al. (2016) and C4 Raffel et al. (2020) datasets, using a context length of 2048 for Llama 1 and 4096 for Llama 2. Results are reported in Table 1. Across all settings, GLVQ consistently achieves lower perplexity than existing methods, particularly in low-bit regimes. For example, under the 2-bit quantization regime on Llama 2-70B, GLVQ-32D achieves a perplexity of 3.36 on Wikitext-2, significantly lower than QTIP (3.78) and QuIP# (3.91), highlighting that the performance gap becomes most pronounced under extreme compression levels. We also observe consistent improvements across model scales and datasets. On Llama 1-13B, GLVQ-32D improves upon QuIP# by 0.41 perplexity points (5.38 vs. 5.79) at 2-bit on Wikitext-2, and 0.53 points (6.95 vs. 7.48) on C4. These results validate the robustness of our method across both small and large Llama variants. Notably, the performance gain is more pronounced as the bit-width decreases, highlighting the advantage of GLVQ’s group-specific lattice codebooks and companding functions. The combination of adaptive vector quantization and nonlinear transformation helps mitigate the quantization error that typically worsens at extreme compression levels. Furthermore, GLVQ-32D consistently outperforms GLVQ-8D, demonstrating that larger lattice dimensions improve quantization fidelity by offering a more expressive codebook structure. This supports our design choice of learning group-specific lattices with adjustable resolution.